# Training LLMs with MXFP4

### v1

@@ Abstract
Low precision (LP) datatypes such as MXFP4 can accelerate matrix multiplications (GEMMs) and reduce training costs. However, directly using MXFP4 instead of BF16 during training significantly degrades model quality. In this work, we present the first near-lossless training recipe that uses MXFP4 GEMMs, which are 2×2$\times$2 × faster than FP8 on supported hardware. Our key insight is to compute unbiased gradient estimates with stochastic rounding (SR), resulting in more accurate model updates. However, directly applying SR to MXFP4 can result in high variance from block-level outliers, harming convergence. To overcome this, we use the random Hadamard tranform to theoretically bound the variance of SR. We train GPT models up to 6.7B parameters and find that our method induces minimal degradation over mixed-precision BF16 training. Our recipe computes $>1/2$ the training FLOPs in MXFP4, enabling an estimated speedup of >1.3×>1.3$\times$> 1.3 × over FP8 and >1.7×>1.7$\times$> 1.7 × over BF16 during backpropagation.
@@ 1 Introduction
Figure 1: Our method uses stochastic rounding (SR) to compute unbiased gradients and the random Hadamard transform to bound the variance of SR. This enables us to perform more accurate model updates with MXFP4 in the backward pass, enabling a speedup of >1.3×>1.3$\times$> 1.3 × over FP8 and >1.7×>1.7$\times$> 1.7 × over BF16.
The latest large language models (LLMs) have billions of parameters that are trained on trillions of tokens, making them incredibly expensive to train. For example, training Llama 3.1 405B required $3\times 10^{24}$ floating point operations (FLOPs), or over 10000 GPUs for multiple months (Dubey et al., 2024). Recent hardware accelerators have started supporting low precision floating point ($\leq$ 16 bit) matrix mulitiplications (GEMMs). Compared to 32-bit GEMMs, hardware-accelerated low precision (LP) GEMMs run at significantly higher throughputs. For example, FP8 GEMMs can be 4×4$\times$4 × faster than FP32 GEMMs and also more energy efficient (NVIDIA, 2024a).
Since LLM training is compute bound in matrix multiplications, LP GEMMs can accelerate training. Almost all modern LLMs are trained with 16 bit GEMMs (Touvron et al., 2023; Dubey et al., 2024), and some even use FP8 GEMMs (Peng et al., 2023). Using 16 bit GEMMs halves the cost of matrix multiplications and improves end-to-end throughput by almost 2×2$\times$2 × (Micikevicius et al., 2018). However, there is no free lunch with low precision training. Reducing the GEMM precision increases quantization distortion and can cause numerical instability.
@@
  • Pretrain GPT models up to 6.7B and show that our training recipe closes the MXFP4-BF16 gap to $<0.1$ validation perplexity.
  • Show that our RHT and SR constructions add minimal overhead to MXFP4 GEMMs, giving a theoretical speedup of >1.3×>1.3$\times$> 1.3 × and >1.7×>1.7$\times$> 1.7 × over a FP8 and BF16 backward pass, respectively.
@@ 2.1 Low Precision Datatypes and IEEE 754 Floating Point Numbers
[…]
@@ 2.3 Mixed Precision Training
One way to accelerate training is with “mixed precision” (MP) training. In MP, parameters are kept in high precision and GEMM operands are converted to a LP datatype for a LP GEMM. MP is a simple way to achieve the throughput benefits of LP datatypes since quantization usually has minimal overhead. End to end, BF16 MP is often $>70\%$ faster than FP32 training (Shoeybi et al., 2020). However, quantization introduces distortion in the GEMM operands and thus outputs. Since the forward and backward passes all happen in low precision, both the loss and the model updates can deviate from their “true” values. At low bitrates $\ll 16$, distortion can degrade model quality and even cause divergence, necessitating advanced training recipes. For example, FP8 MP recipes typically use E4M3 (more precision) in the forward pass and E5M2 (more range) in the backward pass due to the different properties of gradients, weights, and activations (Peng et al., 2023; Engine, ).
At 4 bits, quantization distortion becomes even more difficult to manage. Xi et al. (2023) train smaller non-GPT transformers with INT4 GEMMs by using the non-randomized Hadamard transform in the forward pass and leverage score sampling (LSS) in the backward pass. Since LSS introduces additional overhead, they were only able to achieve an end-to-end speedup of 30% over FP16, which is on par with FP8 mixed precision training (Peng et al., 2023). We are also aware of a concurrent work by Wang et al. (2025) that trains LLMs with FP4. There, the authors train billion parameter GPT models with FP4 in both the forward and backward pass by using a differentiable gradient estimator and keeping outliers in high precision, resulting in a perplexity gap of $>0.5$. Since their work was released after our paper went through the review process, we reserve a full comparison for future work.
@@ 2.4 Stochastic Rounding
$\text{SR}_{\text{dither}}(x)$ =$=\begin{cases}\lfloor x\rfloor&x+\delta<\lfloor x\rfloor+\frac{1}{2}\\ \lceil x\rceil&x+\delta\geq\lfloor x\rfloor+\frac{1}{2}\end{cases}$ (2)
@@ 2.5 Microscaling (MX) FP Formats
0:  $\texttt{emax}_{\texttt{elem}}$ = exponent of largest normal in LP_DTYPE, $k=32$ for hardware support.
@@ 3 Training with MXFP4
Our method hinges on obtaining unbiased, low-variance gradient estimates with pure-MXFP4 GEMMs in the backward pass, enabling more accurate model updates. Since the backward pass consists of $>1/2$ training FLOPs, our recipe can significantly accelerate training without reducing the representational power of the model from LP forward passes (Kumar et al., 2025). To do this, we first modify the OCP MX quantization algorithm to perform unbiased quantization with scaling and stochastic rounding. Then, we show that by first transforming the GEMM operands with a memory-bound construction of the random Hadamard transform (RHT) before quantization, we can bound the variance of the GEMM output. Our method adds minimal overhead while significantly improving the quality of trained models, making MXFP4 practical for training.
@@ 3.1 Unbiased Quantization to MXFP4
Algorithm 1 describes the “reference” MX quantization algorithm to convert a scalar matrix to an MX matrix. Algorithm 1 finds, for each group of 32 entries, value with the largest magnitude $m=\max_{i}(|V_{i}|)$. Then, it calculates a shared exponent as a function of $m$ and $\texttt{emax}_{\texttt{elem}}$, the largest exponent of a normal number in the base data format. For example, $\texttt{emax}_{\texttt{elem}}=2$ for FP4 since its maximum normal value is $1.56=2^{2}*$ 1.5.
@@
[…]
@@
0:  $\texttt{emax}_{\texttt{elem}}$ = exponent of the largest normal number in LP_DTYPE
@@ 3.2 Bounding the Variance of SR with the Random Hadamard Transform
1:  $H\leftarrow\text{Hadamard matrix }H_{b}\in\mathbb{R}^{m\times m}$.
@@
3:  $G^{\prime}\leftarrow\left(\left(\frac{dL}{dy}\right).\texttt{view}\left(\frac{bm}{g},g\right)\right)\texttt{diag}(S)H$
4:  $W^{\prime}\leftarrow H^{T}\texttt{diag}(S)\left(W.\texttt{view}\left(g,\frac{nm}{g}\right)\right)$
5:  $GT^{\prime}\leftarrow\left(\left(\frac{dL}{dy}^{T}\right).\texttt{view}\left(\frac{bm}{g},g\right)\right)\texttt{diag}(S)H$
6:  $X^{\prime}\leftarrow H^{T}\texttt{diag}(S)\left(x.\texttt{view}\left(\frac{bn}{g},g\right)\right)$
7:  $\frac{dL}{dx}\leftarrow\texttt{MXFP4\_GEMM}(G^{\prime},W^{\prime})$
8:  $\frac{dL}{dW}\leftarrow\texttt{MXFP4\_GEMM}(GT^{\prime},X^{\prime})$Where MXFP4_GEMM forms MX groups along the reduction dimension and uses either Algorithm 1 or 2 to quantize to MXFP4.}
@@ 4.1 GPT Pretraining Results
Figure 6: Validation perplexity for training GPT 1.3B for 210 billion tokens, or 5×5$\times$5 × longer than in Table 2. All experiments used BF16 in the forward pass and the specified backward precision.
@@
MXFP4$\star$ Tulu V2 25.9 49.9 62.9 60.5 51.8
Table 3: GPT 6.7B model trained on 20B tokens before and after Tulu V2 fine-tuning. Both BF16 and our MXFP4+RHT+SR (MXFP4$\star$) model exhibit similar performance before and after fine-tuning.
@@
Table 5: Throughput for a FP16 forward pass and specified backward pass of a Llama 2 70B decoder layer. Measured on a NVIDIA A100; see Section 4.2 for more details. Since the A100 can perform INT4 GEMMs 4×4$\times$4 × faster than FP16 GEMMs, these numbers represent the expected speedup of MXFP4 on supported hardware.
@@ 4.2 Overhead Calculations
Second, we measured the overhead of the RHT on the HuggingFace implementation (Wolf et al., 2020) of single Llama 2 70B decoder layer decoder layer on a NVIDIA A100 GPU. Specifically, we report the end-to-end tokens per second for computing the forward pass in FP16 and backward pass in either FP16 or INT4, which has the same hardware speedup (4×4$\times$4 ×) on the NVIDIA A100 vs. FP16 as MXFP4 has on modern hardware. We also include INT8 as a proxy for the expected speedup of a FP8 backward pass. We use a batch size of 4 sequences with 4K tokens each (16K tokens/batch), Flash Attention 2, torch.compile, and the CUTLASS INT4 and INT8 GEMM kernels. We were unable to use CUDA graphs since the HuggingFace implementation is not compatible with CUDA graphs; we expect CUDA graphs to improve speedup ratios by masking kernel launch overhead.
@@
To measure the overhead of stochastic rounding, we used an Amazon Trainium 1 chip (EC2 Trn1 instance), which is one of the few widely available chips that has dedicated stochastic rounding hardware (Amazon, ). Our experiments show that for most matrix sizes, using SR to quantize GEMM operands from FP32 to BF16 adds less than 2% overhead over the BF16 GEMM itself. Assuming a 4$\times$ increase in GEMM throughput when going from BF16 to FP4, this would mean SR adds less than 10% overhead.
@@ 6.1 FP8 Forward Pass Results
This section contains experiments using mixed-precision FP8 in the forward pass and MXFP4 in the backward pass. Prior works have shown that mixed-precision FP8 forward and backward passes can be close to lossless over mixed-precision BF16 training (Peng et al., 2023; NVIDIA, b). To test if MXFP4 backward passes are still practical with FP8 forward passes, we trained GPT 1.3B and 6.7B models with NVIDIA’s TransformerEngine (TE) FP8 (E4M3) implementation (NVIDIA, b) in the forward pass and our MXFP4 formulation in the backward pass. We did not test GPT 345M since TE FP8 already had a $>0.1$ validation perplexity gap vs. BF16 in our experiments. For GPT 6.7B, since we did not have access to FP8-capable hardware with fast interconnects necessary for tensor-parallel training, we emulated FP8 matrix multiplications by dequantizing FP8 GEMM operands into BF16 and performing a BF16 GEMM. While this is not bit-accurate vs. a FP8 GEMM, the relative error of the output is $\approx 0.3\%$ for random Gaussian inputs. Furthermore, this is essentially how PyTorch emulates FP8 GEMMs (Team, ). At both model scales, we find that FP8 forward passes and MXFP4 backward passes are sufficient to essentially match BF16 training.
@@ 7 Experimental Setup Details
All experiments were run on AWS P4 and G6e EC2 instances. Our code was based off of the Megatron-LM codebase at Github commit a4ad305d4b117217141730b9b18af52dda069450 and the Microsoft microxcaling codebase at Github commit 7bc41952de394f5cc5e782baf132e7c7542eb4e4. We used the NVIDIA Pytorch + Ubuntu 24.04 docker image, which contains a version of Transformer Engine 1.5 for the FP8 experiments. All models were trained with the AdamW optimizer, FlashAttention, and the following hyperparameters:
@@
[…]

### v3

@@ Abstract
Low precision (LP) datatypes such as MXFP4 can accelerate matrix multiplications (GEMMs) and reduce training costs. However, directly using MXFP4 instead of BF16 during training significantly degrades model quality. In this work, we present the first near-lossless training recipe that uses MXFP4 GEMMs, which are $2\times$ faster than FP8 on supported hardware. Our key insight is to compute unbiased gradient estimates with stochastic rounding (SR), resulting in more accurate model updates. However, directly applying SR to MXFP4 can result in high variance from block-level outliers, harming convergence. To overcome this, we use the random Hadamard tranform to theoretically bound the variance of SR. We train GPT models up to 6.7B parameters and find that our method induces minimal degradation over mixed-precision BF16 training. Our recipe computes >1/2>1/2 the training FLOPs in MXFP4, enabling an estimated speedup of $1.3\times$ over FP8 and $1.7\times$ over BF16 during backpropagation.
@@ 1 Introduction
Figure 1: Our method uses stochastic rounding (SR) to compute unbiased gradients and the random Hadamard transform to bound the variance of SR. This enables us to perform more accurate model updates with MXFP4 in the backward pass, enabling a speedup of $1.3\times$ over FP8 and $1.7\times$ over BF16.
The latest large language models (LLMs) have billions of parameters that are trained on trillions of tokens, making them incredibly expensive to train. For example, training Llama 3.1 405B required $3\times 10^{24}$ floating point operations (FLOPs), or over 10000 GPUs for multiple months (Dubey et al., 2024). Recent hardware accelerators have started supporting low precision floating point ($\leq$ 16 bit) matrix mulitiplications (GEMMs). Compared to 32-bit GEMMs, hardware-accelerated low precision (LP) GEMMs run at significantly higher throughputs. For example, FP8 GEMMs can be $4\times$ faster than FP32 GEMMs and also more energy efficient (NVIDIA, 2024a).
Since LLM training is compute bound in matrix multiplications, LP GEMMs can accelerate training. Almost all modern LLMs are trained with 16 bit GEMMs (Touvron et al., 2023; Dubey et al., 2024), and some even use FP8 GEMMs (Peng et al., 2023). Using 16 bit GEMMs halves the cost of matrix multiplications and improves end-to-end throughput by almost $2\times$ (Micikevicius et al., 2018). However, there is no free lunch with low precision training. Reducing the GEMM precision increases quantization distortion and can cause numerical instability.
@@
  • Pretrain GPT models up to 6.7B and show that our training recipe closes the MXFP4-BF16 gap to <0.1<0.1 validation perplexity.
  • Show that our RHT and SR constructions add minimal overhead to MXFP4 GEMMs, giving a theoretical speedup of $1.3\times$ and $1.7\times$ over a FP8 and BF16 backward pass, respectively.
@@ 2.1 Low Precision Datatypes and IEEE 754 Floating Point Numbers
IEEE floats (Table 1) are defined with 1 sign bit, $e$ exponent bits, and $m$ mantissa bits. In shorthand, a $1+m+e$ bit datatype is written as EeeMmm. The actual “normal” value represented by an IEEE float with sign bit $S$, mantissa $M$, and exponent $E$ is
@@ 2.3 Mixed Precision Training
One way to accelerate training is with “mixed precision” (MP) training. In MP, parameters are kept in high precision and GEMM operands are converted to a LP datatype for a LP GEMM. MP is a simple way to achieve the throughput benefits of LP datatypes since quantization usually has minimal overhead. End to end, BF16 MP is often $70\%$ faster than FP32 training (Shoeybi et al., 2020). However, quantization introduces distortion in the GEMM operands and thus outputs. Since the forward and backward passes all happen in low precision, both the loss and the model updates can deviate from their “true” values. At low bitrates $\ll 16$, distortion can degrade model quality and even cause divergence, necessitating advanced training recipes. For example, FP8 MP recipes typically use E4M3 (more precision) in the forward pass and E5M2 (more range) in the backward pass due to the different properties of gradients, weights, and activations (Peng et al., 2023; Engine, ).
At 4 bits, quantization distortion becomes even more difficult to manage. Xi et al. (2023) train smaller non-GPT transformers with INT4 GEMMs by using the non-randomized Hadamard transform in the forward pass and leverage score sampling (LSS) in the backward pass. Since LSS introduces additional overhead, they were only able to achieve an end-to-end speedup of 30% over FP16, which is on par with FP8 mixed precision training (Peng et al., 2023). We are also aware of a concurrent work by Wang et al. (2025) that trains LLMs with FP4. There, the authors train billion parameter GPT models with FP4 in both the forward and backward pass by using a differentiable gradient estimator and keeping outliers in high precision, resulting in a perplexity gap of >0.5>0.5. Since their work was released after our paper went through the review process, we reserve a full comparison for future work.
@@ 2.4 Stochastic Rounding
$\text{SR}_{\text{dither}}(x)$ $=\begin{cases}\lfloor x\rfloor&x+\delta<\lfloor x\rfloor+\frac{1}{2}\\ \lceil x\rceil&x+\delta\geq\lfloor x\rfloor+\frac{1}{2}\end{cases}$ (2)
@@ 2.5 Microscaling (MX) FP Formats
[…]
@@ 3 Training with MXFP4
Our method hinges on obtaining unbiased, low-variance gradient estimates with pure-MXFP4 GEMMs in the backward pass, enabling more accurate model updates. Since the backward pass consists of >1/2>1/2 training FLOPs, our recipe can significantly accelerate training without reducing the representational power of the model from LP forward passes (Kumar et al., 2025). To do this, we first modify the OCP MX quantization algorithm to perform unbiased quantization with scaling and stochastic rounding. Then, we show that by first transforming the GEMM operands with a memory-bound construction of the random Hadamard transform (RHT) before quantization, we can bound the variance of the GEMM output. Our method adds minimal overhead while significantly improving the quality of trained models, making MXFP4 practical for training.
@@ 3.1 Unbiased Quantization to MXFP4
Algorithm 1 describes the “reference” MX quantization algorithm to convert a scalar matrix to an MX matrix. Algorithm 1 finds, for each group of 32 entries, value with the largest magnitude $m=\max_{i}(|V_{i}|)$. Then, it calculates a shared exponent as a function of $m$ and $\texttt{emax}_{\texttt{elem}}$, the largest exponent of a normal number in the base data format. For example, $\texttt{emax}_{\texttt{elem}}=2$ for FP4 since its maximum normal value is $6=2^{2}*1.5$.
@@
For MXFP4, line 1 of Algorithm 1 returns $\texttt{shared\_exp}\leftarrow\lfloor\log_{2}(m)\rfloor-2$. Observe that after dividing the entire group by $2^{\texttt{shared\_exp}}$, $m$ becomes
@@
[…]
@@ 3.2 Bounding the Variance of SR with the Random Hadamard Transform
[…]
@@
[…]
@@ 4.1 GPT Pretraining Results
Figure 6: Validation perplexity for training GPT 1.3B for 210 billion tokens, or $5\times$ longer than in Table 2. All experiments used BF16 in the forward pass and the specified backward precision.
@@
[…]
@@
Table 5: Throughput for a FP16 forward pass and specified backward pass of a Llama 2 70B decoder layer. Measured on a NVIDIA A100; see Section 4.2 for more details. Since the A100 can perform INT4 GEMMs $4\times$ faster than FP16 GEMMs, these numbers represent the expected speedup of MXFP4 on supported hardware.
@@ 4.2 Overhead Calculations
Second, we measured the overhead of the RHT on the HuggingFace implementation (Wolf et al., 2020) of single Llama 2 70B decoder layer decoder layer on a NVIDIA A100 GPU. Specifically, we report the end-to-end tokens per second for computing the forward pass in FP16 and backward pass in either FP16 or INT4, which has the same hardware speedup ($4\times$) on the NVIDIA A100 vs. FP16 as MXFP4 has on modern hardware. We also include INT8 as a proxy for the expected speedup of a FP8 backward pass. We use a batch size of 4 sequences with 4K tokens each (16K tokens/batch), Flash Attention 2 (Dao, 2024), torch.compile, and the CUTLASS INT4 and INT8 GEMM kernels. We were unable to use CUDA graphs since the HuggingFace implementation is not compatible with CUDA graphs; we expect CUDA graphs to improve speedup ratios by masking kernel launch overhead.
@@
[…]
@@ 6.1 FP8 Forward Pass Results
This section contains experiments using mixed-precision FP8 in the forward pass and MXFP4 in the backward pass. Prior works have shown that mixed-precision FP8 forward and backward passes can be close to lossless over mixed-precision BF16 training (Peng et al., 2023; NVIDIA, b). To test if MXFP4 backward passes are still practical with FP8 forward passes, we trained GPT 1.3B and 6.7B models with NVIDIA’s TransformerEngine (TE) FP8 (E4M3) implementation (NVIDIA, b) in the forward pass and our MXFP4 formulation in the backward pass. We did not test GPT 345M since TE FP8 already had a >0.1>0.1 validation perplexity gap vs. BF16 in our experiments. For GPT 6.7B, since we did not have access to FP8-capable hardware with fast interconnects necessary for tensor-parallel training, we emulated FP8 matrix multiplications by dequantizing FP8 GEMM operands into BF16 and performing a BF16 GEMM. While this is not bit-accurate vs. a FP8 GEMM, the relative error of the output is $\approx 0.3\%$ for random Gaussian inputs. Furthermore, this is essentially how PyTorch emulates FP8 GEMMs (Team, ). At both model scales, we find that FP8 forward passes and MXFP4 backward passes are sufficient to essentially match BF16 training.
@@ 7 Experimental Setup Details
All experiments were run on AWS P4 and G6e EC2 instances. Our code was based off of the Megatron-LM codebase at Github commit a4ad305d4b117217141730b9b18af52dda069450 and the Microsoft microxcaling codebase at Github commit 7bc41952de394f5cc5e782baf132e7c7542eb4e4. We used the NVIDIA Pytorch + Ubuntu 24.04 docker image, which contains a version of Transformer Engine 1.5 for the FP8 experiments. All models were trained with the AdamW optimizer, FlashAttention (Dao et al., 2022), and the following hyperparameters:
@@
By definition, $\texttt{stochastic\_round\_to\_FP4}(x)$ produces an unbiased estimate of $x$ as long as $x$ is “within range” – i.e. it does not overflow outside of the range of representable values in FP4. Since the maximum normal in FP4 is 6, $\texttt{stochastic\_round\_to\_FP4}(w)$ will give an unbiased FP4 estimate of $\frac{3}{4}v/X$. Finally, from linearity of expectation, $X*\texttt{stochastic\_round\_to\_FP4}(w)$ gives an unbiased estimate of $\frac{3}{4}v$, as desired.